OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2

OpenStudy (nurali):

-4(2x - 1) = -6(x + 2) - 2 -8x+4=-6x-12-2 -8x+6x=-14-4 -2x=-18 x=9

OpenStudy (dobby1):

darn you beat me to it

OpenStudy (dobby1):

good job

OpenStudy (anonymous):

Ohhh, okay. I see what I've done wrong. Thank you!

OpenStudy (nurali):

Welcome to Openstudy

OpenStudy (nurali):

My Pleasure